Graph the two equations to solve the system. What is the ordered pair solution?
x + y = 6 and x - y = 4

(4, 2)

(6, 2)

(5, 1)

no solution

Answer by mananth(16949)

1 x + 1 y = 6 .............1
1 x -1 y = 4 .............2
equation1
when x= 0 y= 6 ( 0 , 6 )
when y = 0 x= 6 ( 6 , 0 )
Equation 2
when x= 0 y= -4 ( 0 , -4 )
when y = 0 x= 4 ( 4 , 0 )


(5, 1)